A solution containing 4.5 g of glycerol, a nonvolatile nonelectrolyte, in 100.0 g of ethanol has a boiling point of 79.0oC. If the normal BP of ethanol is 78.4oC, calculate the molar mass of glycerol. Given: DTb = mass solute = mass solvent = Kb = 1.22oC/m (Table 13.4)
Find: molar mass (g/mol) DTb = Kb  m
1
Expert's answer
2018-02-24T07:58:08-0500
DTb = Kb ∙ m
79.0 – 78.4 = 1.22 ∙ m
m = 0.6 / 1.22 ≈ 0.49
m = 4.5 g / M (glycerol) : 0.1 kg
So
4.5 / M (glycerol) : 0.1 = 0.49
M (glycerol) = 91.5
Answer: 91.5 g/mol.